# Functional Distribution Networks (FDN)

### v2

@@ 1 Introduction
Modern neural predictors are routinely deployed under dataset shift, where test inputs depart from the training distribution. In these regimes, point predictions from deterministic networks and naïve uncertainty surrogates from traditional stochastic heuristics often become overconfident—assigning high probability to wrong outcomes, particularly off-support/extrapolation—undermining reliable decision making. Bayesian Neural Networks (BNNs), MLP with dropout, Deep Ensembles, and Hypernetworks are strong practical baselines, yet they can still under-react outside the training support or require substantial ensembling/sampling to behave robustly (quinonero2009dataset; ovadia2019can; guo2017calibration; mackay1992practical; neal1996bayesian; graves2011practical; blundell2015weight; gal2016dropout; lakshminarayanan2017simple; ha2017hypernetworks).
This motivates architectures that are uncertainty-aware and calibrated—sharp in-distribution (ID) and widening appropriately OOD (higher CRPS, wider intervals). We pursue this with Functional Distribution Networks (FDN), which place input-conditional distributions over weights to modulate uncertainty locally in $x$. Concretely, FDN amortizes an input-conditional posterior $q_{\phi}(\theta\mid x)$ via small Hypernetworks and trains it with a Monte Carlo likelihood and a $\beta$-ELBO. Under matched budgets, FDN is ID-competitive and well-calibrated by scale on controlled 1D shifts: on low-frequency or piecewise-smooth tasks (step, quadratic) it achieves near-unity MSE–Var slopes with high Spearman correlation and favorable AURC/CRPS, indicating that predictive variance tracks error growth both in- and out-of-distribution. On more oscillatory shifts (sine) FDN preserves strong rank calibration and increases variance OOD, but its variance still under-scales relative to the error—an explicit target for future improvement. Beyond 1D toy functions, we evaluate We further evaluate FDN on standard small/medium UCI-style regression benchmarks—Airfoil Self-Noise, Combined Cycle Power Plant, and Energy Efficiency (uci_airfoil; uci_ccpp; uci_energy)—under feature-based ID/OOD splits. On these real datasets FDN remains ID-competitive and typically exhibits large positive $\Delta\mathrm{Var}$ with moderate $\Delta\mathrm{MSE}$ and $\Delta\mathrm{CRPS}$, i.e., it widens uncertainty under shift while maintaining sensible accuracy and calibration relative to strong Bayesian, ensemble, and hypernetwork baselines.
@@
Contributions.
(i) Model. We introduce Functional Distribution Networks (FDN), a simple module that amortizes input-conditioned weight distributions via small Hypernetworks, in two variants: IC-FDNet (conditioning on $x$) and LP-FDNet (conditioning layer-wise on previous activations). (ii) Protocol. We propose a small-suite extrapolation protocol that targets $\Delta\mathrm{Var}>0$ under shift and complements it with calibration diagnostics (MSE–variance slope, rank correlation, and risk–coverage). (iii) Empirical study. Under matched parameter, update, and predictive-sample budgets, we benchmark FDN against strong Bayesian, ensemble, dropout, and hypernetwork baselines on controlled 1D function families and UCI-style regression tasks, showing that FDN remains ID-competitive while providing practically useful, shift-aware uncertainty.
Scope.
We focus on low-dimensional regression with homoscedastic scalar Gaussian heads and relatively shallow backbones in order to isolate ID vs. OOD behavior under tightly matched budgets. All code, configurations, and scripts to reproduce the experiments will be released upon acceptance.
@@ 2 Related Work
Uncertainty in neural regression. BNN place distributions over weights and infer posteriors via variational approximations or MCMC (mackay1992practical; neal1996bayesian; graves2011practical; blundell2015weight). Deep Ensembles average predictions from independently trained networks and are a strong practical baseline (lakshminarayanan2017simple; maddox2019simple). MLP (with dropout) interprets dropout at test time as approximate Bayesian inference (gal2016dropout). Heteroscedastic regression learns input-dependent output variance but retains deterministic weights (nix1994estimating; kendall2017uncertainties).
Hypernetworks and conditional weight generation. Hypernetworks generate the weights of a primary network using an auxiliary network (ha2017hypernetworks); related work explores dynamic, input-conditioned filters and conditional computation (debrabandere2016dynamic; brock2018smash). Bayesian/uncertainty-aware Hypernetworks place distributions over generated weights and train them variationally (krueger2017bayesian). Our FDN differs by explicitly conditioning the weight distributions on the current input (or intermediate activations) in order to modulate epistemic uncertainty itself, not only deterministic weights. This input-aware weight stochasticity is what enables FDN to widen uncertainty under distributional shift while maintaining competitive ID accuracy under matched budgets.
Meta-learning and context-conditioned predictors. Neural Processes (NP) learn a distribution over functions conditioned on a context set $\mathcal{C}$ via a global latent, yielding $p(y\mid x,\mathcal{C})$ (garnelo2018neural; kim2019attentive); gradient-based meta-learning (e.g., MAML) instead adapts an initialization per task via inner-loop gradients (finn2017model). FDN is complementary: it amortizes local weight uncertainty directly on $x$ (and optionally a compact context) through $q_{\phi}(\theta\mid x,c)$, requires no inner-loop adaptation, and injects uncertainty at the layer level rather than via a single global latent.
Calibration and OOD behavior. Proper scoring rules such as the continuous ranked probability score (CRPS) are strictly proper and reward calibrated predictive distributions (gneiting2007strictly). Empirical studies highlight overconfidence under dataset shift and introduce OOD benchmarks (ovadia2019can). Our evaluation therefore separates interpolation from extrapolation and uses the monotonic relationship between per-sample squared error (MSE) and predicted variance as a simple diagnostic calibration check.
@@ 3.1 Preliminaries
We model $y\in\mathbb{R}^{d_{y}}$ with a neural network $f_{\theta}:\mathbb{R}^{d_{x}}\!\to\!\mathbb{R}^{D_{y}}$ ($D_{y}=2d_{y}$ for a diagonal covariance and $D_{y}=d_{y}(d_{y}+1)/2$ for full covariance) whose final layer parameterizes a Gaussian predictive head,
$f_{\theta}(x)=\big(\mu_{\theta}(x),\Sigma_{\theta}(x)\big),\qquad p(y\mid x,\theta)=\mathcal{N}\!\big(y;\,\mu_{\theta}(x),\,\Sigma_{\theta}(x)\big),$
In the main experiments we restrict to scalar outputs and use a homoscedastic Gaussian likelihood $p(y\mid x,\theta)=\mathcal{N}\!\bigl(y;\mu_{\theta}(x),\sigma^{2}\bigr)$ with fixed variance $\sigma^{2}$, shared across all models. For such a head the per-example contribution to the $\beta$–ELBO reduces to a rescaled squared-error term plus the KL penalty between the input-conditioned weight posterior and the prior; constants can be absorbed into the trade-off parameter $\beta$. We therefore optimize a weighted $MSE+\beta KL$ objective and use the same Gaussian head for all baselines to ensure a fair comparison. More general multivariate, heteroscedastic, and structured-covariance likelihoods are discussed in Appendix D.
@@ 3.2 FDN: Input-Conditioned Weight Distributions
$(\mu_{W,\ell},\,\rho_{W,\ell},\,\mu_{b,\ell},\,\rho_{b,\ell})\;=\;A_{\ell}(s_{\ell}^{(k)}),\qquad\sigma_{W,\ell}\;=\;\varepsilon+\operatorname{softplus}(\rho_{W,\ell}),\quad\sigma_{b,\ell}\;=\;\varepsilon+\operatorname{softplus}(\rho_{b,\ell}),$
with a small floor $\varepsilon=10^{-3}$ for numerical stability. Sampling then proceeds as
$W_{\ell}^{(k)}=\mu_{W,\ell}+\sigma_{W,\ell}\odot z_{W,\ell}^{(k)},\qquad b_{\ell}^{(k)}=\mu_{b,\ell}+\sigma_{b,\ell}\odot z_{b,\ell}^{(k)},\qquad z_{\{\cdot\},\ell}^{(k)}\sim\mathcal{N}(0,I),$
with the conditioning signal chosen as $s_{\ell}=x$ for IC-FDNet and $s_{\ell}=a_{\ell-1}^{(k)}$ for LP-FDNet, where in both cases the layer output is $a_{\ell}^{(k)}=f_{\ell}(a_{\ell-1}^{(k)};W_{\ell}^{(k)},b_{\ell}^{(k)})$ (sequential across layers). Figure 1 illustrates a single FDN layer: a small hypernetwork maps $s_{\ell}$ to the parameters of a Gaussian distribution over weights and biases; a sample $\theta_{\ell}=(W_{\ell},b_{\ell})$ from this distribution is then used to compute $a_{\ell}=f_{\ell}(a_{\ell-1};\theta_{\ell})$.
@@
$q_{\phi}(\theta\mid x)=\prod_{\ell=1}^{L}\mathcal{N}\!\big(\mathrm{vec}(W_{\ell});\mu_{W,\ell}(s_{\ell}),\,\operatorname{diag}\sigma_{W,\ell}^{2}(s_{\ell})\big)\,\mathcal{N}\!\big(b_{\ell};\mu_{b,\ell}(s_{\ell}),\,\operatorname{diag}\sigma_{b,\ell}^{2}(s_{\ell})\big),$
@@
$q_{\phi}(\theta\mid x)=\mathcal{N}\!\big(\theta;\,\mu_{\phi}(x),\,\operatorname{diag}\sigma_{\phi}^{2}(x)\big),\qquad\theta^{(k)}=\mu_{\phi}(x)+\sigma_{\phi}(x)\odot\varepsilon^{(k)},\ \ \varepsilon^{(k)}\sim\mathcal{N}(0,I),$
@@
$D_{\mathrm{KL}}\!\left(\mathcal{N}(\mu,\operatorname{diag}\sigma^{2})\,\middle\|\,\mathcal{N}(0,\sigma_{0}^{2}I)\right)=\tfrac{1}{2}\sum_{j}\!\left(\frac{\sigma_{j}^{2}+\mu_{j}^{2}}{\sigma_{0}^{2}}\;-\;1\;-\;\log\frac{\sigma_{j}^{2}}{\sigma_{0}^{2}}\right),$
and we sum this over all layers (for both $W_{\ell}$ and $b_{\ell}$). The variance floor is implemented by the $\varepsilon$ in $\sigma=\varepsilon+\mathrm{softplus}(\rho)$ rather than a hard bound.
@@ 4 Experiments
We evaluate FDN on (i) controlled 1D regression tasks where the ground-truth function is known and interpolation vs. extrapolation is precisely defined in input space, and (ii) standard small/medium UCI-style regression benchmarks with feature-based ID/OOD splits. In both settings we report metrics separately on the interpolation (ID) region, the extrapolation (OOD) region, and the aggregated test set, and we summarize distribution shift by deltas $\Delta(\cdot)=\mathbb{E}_{\mathrm{OOD}}[\cdot]-\mathbb{E}_{\mathrm{ID}}[\cdot]$.
@@ 4.3 Complexity, capacity, and fairness
Link-budget.
We consider networks with a single hidden layer and fix the parameter budget to $P\approx 1000$ (±5%) for all models, counting all trainable parameters, including any Hypernetwork components; counts appear in Table 4 (Appendix F). To equalize the update budget, ensembles with $M$ members use epoch-split training (epochs divided by $M$). For non-ensemble networks we use one Monte Carlo draw per update ($K=1$), keeping per-step cost comparable and the total number of parameter updates matched across models.
@@ 4.4 Metrics and calibration diagnostics
To avoid cherry-picking individual runs, all qualitative plots are based on a representative seed chosen by a fixed aggregation procedure. For each dataset and configuration we run all models over multiple random seeds (20 seeds per toy task, 100 for Airfoil, and 3 for the remaining UCI datasets) and collect summary metrics per seed (e.g., $\text{MSE}_{\text{ID}}$, $\text{MSE}_{\text{OOD}}$, $\text{Var}_{\text{ID}}$, $\text{Var}_{\text{OOD}}$, $\Delta\text{CRPS}$, AURC). For IC-FDNet (or LP-FDNet) we compute the coordinate-wise median of these metrics over seeds and then select the seed whose metric vector is closest (in Euclidean distance) to this median. All per-seed plots in the main text are generated from this representative seed, using the same rule for every dataset, so that the shown behavior is typical rather than hand-picked. On the Airfoil Self-Noise dataset, Figure 5 (Appendix F) visualizes the resulting seed-aggregated MSE–variance scatter, confirming that the selected representative seed lies close to the overall across-seed trend.
@@ 4.5 Results
Real regression benchmarks.
On the real regression tasks (Airfoil, CCPP, Energy), the same patterns largely persist (Figures 3 and 4; Tables 810). FDN achieves reasonable in-distribution MSE and typically exhibits large positive $\Delta\mathrm{Var}$, indicating that uncertainty widens under feature-based shift. This comes with somewhat larger $\Delta\mathrm{MSE}$ and $\Delta\mathrm{CRPS}$ than the sharpest baselines on Airfoil, but more moderate values on CCPP and Energy, so calibration degrades gradually rather than catastrophically. On Airfoil, for example, FDN’s ID scatter lies close to the ideal MSE==Var line and spreads out smoothly OOD, with strong Spearman correlation; on CCPP and Energy the predictive variances remain shift-aware and provide useful selective-risk behavior, even when the absolute scale is not always better than the strongest baselines. Taken together, the toy and UCI results suggest that input-conditioned weight stochasticity is a viable and modular route to OOD-aware regression: FDN behaves like a drop-in “uncertainty layer” that can be tuned via $\beta$ and hypernetwork capacity to trade off in-distribution sharpness against OOD conservatism.
@@ 5 Limitations
FDN’s input-conditioned weight stochasticity, like similar stochastic layers, can overfit spurious cues if $\beta$ is too small or the prior is too loose; careful KL scheduling and priors are important. LP-FDN samples weights layer-by-layer, adding latency versus a deterministic pass; sampling at test time also incurs a compute/latency trade-off with $K$, although in our experiments we keep $K$ small. Our study focuses on low-dimensional regression with small/medium tabular datasets and relatively shallow architectures; scaling to high-dimensional inputs (e.g., images), deeper backbones, and structured outputs will require additional engineering, such as low-rank or adapter-style Hypernetworks and more aggressive capacity control.
On the calibration side, FDN still under-scales variance on highly oscillatory OOD regimes (the sine task), and nothing in the current design explicitly enforces frequency-aware or spectral robustness. We also restrict attention to homoscedastic scalar Gaussian likelihoods; heteroscedastic and full-covariance heads, as well as structured priors that tie together different layers or groups of weights, are left to future work. Finally, we only consider regression; applying FDN to classification would require discrete predictive mixtures and calibration metrics beyond CRPS (e.g., ECE/Brier), and may interact non-trivially with common tricks such as label smoothing or temperature scaling.

### v3

@@ 1 Introduction
Modern neural predictors are often deployed under dataset shift, where test inputs depart from the training distribution. In such regimes, deterministic networks and common stochastic heuristics frequently become overconfident, assigning high probability to incorrect predictions far from the training support and undermining reliable decision making. Bayesian Neural Networks (BNNs), MLPs with dropout, Deep Ensembles, and Hypernetworks are strong practical baselines, yet they can still under-react under shift or require substantial ensembling or sampling to achieve robust uncertainty behavior (Quiñonero-Candela et al., 2009; Ovadia et al., 2019; Guo et al., 2017; MacKay, 1992; Neal, 1996; Graves, 2011; Blundell et al., 2015; Gal and Ghahramani, 2016; Lakshminarayanan et al., 2017; Ha et al., 2017).
This motivates architectures that are explicitly uncertainty-aware and calibrated, producing sharp predictions in-distribution (ID) while widening appropriately under distribution shift. We pursue this goal with Functional Distribution Networks (FDN), which place input-conditioned distributions over network weights, allowing predictive uncertainty to adapt locally in input space. FDN amortizes an input-conditional posterior $q_{\phi}(\theta\mid x)$ via small hypernetworks and is trained using a Monte Carlo likelihood and a $\beta$-ELBO objective.
Empirically, under matched parameter and inference budgets, FDN is ID-competitive and exhibits strong calibration under shift. On controlled 1D regression tasks, predictive variance closely tracks error growth in- and out-of-distribution, while on more challenging oscillatory shifts FDN preserves rank calibration and increases uncertainty OOD, albeit with some under-scaling. We further evaluate FDN on standard UCI-style regression benchmarks—Airfoil Self-Noise, Combined Cycle Power Plant, and Energy Efficiency (Brooks et al., 1989; Tüfekci and Kaya, 2014; Tsanas and Xifara, 2012)—where it typically widens uncertainty under feature-based shifts while maintaining competitive accuracy and calibration relative to strong Bayesian, ensemble, and hypernetwork baselines.
@@
Contributions. (i) Model. We introduce Functional Distribution Networks (FDN), a simple module that amortizes input-conditioned weight distributions via small Hypernetworks, in two variants: IC-FDNet (conditioning on $x$) and LP-FDNet (conditioning layer-wise on previous activations). (ii) Protocol. We propose a small-suite extrapolation protocol that targets $\Delta\mathrm{Var}>0$ under shift and complements it with calibration diagnostics (MSE–variance slope, rank correlation, and risk–coverage). (iii) Empirical study. Under matched parameter, update, and predictive-sample budgets, we benchmark FDN against strong Bayesian, ensemble, dropout, and hypernetwork baselines on controlled 1D function families and UCI-style regression tasks, showing that FDN remains ID-competitive while providing practically useful, shift-aware uncertainty.
Scope. We focus on low-dimensional regression with homoscedastic scalar Gaussian heads and relatively shallow backbones in order to isolate ID vs. OOD behavior under tightly matched budgets. All code, configurations, and scripts to reproduce the experiments will be released upon acceptance.
@@ 2 Related Work
Uncertainty in neural regression. BNN place distributions over weights and infer posteriors via variational approximations or MCMC (MacKay, 1992; Neal, 1996; Graves, 2011; Blundell et al., 2015). Deep Ensembles average predictions from independently trained networks and are a strong practical baseline (Lakshminarayanan et al., 2017; Maddox et al., 2019). MLP (with dropout) interprets dropout at test time as approximate Bayesian inference (Gal and Ghahramani, 2016). Heteroscedastic regression learns input-dependent output variance but retains deterministic weights (Nix and Weigend, 1994; Kendall and Gal, 2017).
Hypernetworks and conditional weight generation. Hypernetworks generate the weights of a primary network using an auxiliary network (Ha et al., 2017); related work explores dynamic, input-conditioned filters and conditional computation (De Brabandere et al., 2016; Brock et al., 2018). Bayesian/uncertainty-aware Hypernetworks place distributions over generated weights and train them variationally (Krueger et al., 2017). Our FDN differs by explicitly conditioning the weight distributions on the current input (or intermediate activations) in order to modulate epistemic uncertainty itself, not only deterministic weights. This input-aware weight stochasticity is what enables FDN to widen uncertainty under distributional shift while maintaining competitive ID accuracy under matched budgets.
Calibration and OOD behavior. Proper scoring rules such as the continuous ranked probability score (CRPS) are strictly proper and reward calibrated predictive distributions (Gneiting and Raftery, 2007). Empirical studies highlight overconfidence under dataset shift and introduce OOD benchmarks (Ovadia et al., 2019). Our evaluation therefore separates interpolation from extrapolation and uses the monotonic relationship between per-sample squared error (MSE) and predicted variance as a simple diagnostic calibration check.
@@ 3.1 Preliminaries
We model $y\in\mathbb{R}^{d_{y}}$ with a neural network $f_{\theta}:\mathbb{R}^{d_{x}}\!\to\!\mathbb{R}^{D_{y}}$ whose final layer parametrizes a Gaussian predictive head $(\mu_{\theta}(x),\Sigma_{\theta}(x))$, yielding
$p(y\mid x,\theta)=\mathcal{N}\!\big(y;\mu_{\theta}(x),\Sigma_{\theta}(x)\big).$
In the main experiments we use a homoscedastic Gaussian likelihood $p(y\mid x,\theta)=\mathcal{N}\!\bigl(y;\mu_{\theta}(x),\sigma^{2}\bigr)$ with fixed variance $\sigma^{2}$, shared across all models. This removes aleatoric uncertainty and isolates epistemic uncertainty arising from stochastic weights. Under this likelihood, the $\beta$–ELBO reduces (up to constants) to a weighted squared-error term plus a KL penalty, yielding a $MSE+\beta\,KL$ objective. We use the same Gaussian head for all baselines to ensure a fair comparison.
More general likelihoods are discussed in Appendix D.
@@ 3.2 FDN: Input-Conditioned Weight Distributions
$(\mu_{W,\ell},\,\rho_{W,\ell},\,\mu_{b,\ell},\,\rho_{b,\ell})$ $=A_{\ell}\!\left(s_{\ell}^{(k)}\right),$
where $\sigma_{.,\ell}=\varepsilon+\operatorname{softplus}(\rho_{.,\ell})$. with a small variance floor $\varepsilon=10^{-3}$ for numerical stability. Sampling then proceeds via the standard Gaussian reparameterization trick. The conditioning signal is chosen as $s_{\ell}=x$ for IC-FDNet and $s_{\ell}=a_{\ell-1}^{(k)}$ for LP-FDNet, where in both cases the layer output is $a_{\ell}^{(k)}=f_{\ell}(a_{\ell-1}^{(k)};W_{\ell}^{(k)},b_{\ell}^{(k)})$ (sequential across layers). Figure 1 illustrates a single FDN layer: a small hypernetwork maps $s_{\ell}$ to the parameters of a Gaussian distribution over weights and biases; a sample $\theta_{\ell}=(W_{\ell},b_{\ell})$ from this distribution is then used to compute $a_{\ell}=f_{\ell}(a_{\ell-1};\theta_{\ell})$.
@@
$q_{\phi}(\theta\mid x)=\prod_{\ell=1}^{L}\mathcal{N}\!\big((W_{\ell},b_{\ell});(\mu_{W,\ell},\mu_{b,\ell}),\operatorname{diag}(\sigma_{W,\ell}^{2},\sigma_{b,\ell}^{2})\big).$
@@
$q_{\phi}(\theta\mid x)=\mathcal{N}\!\big(\theta;\,\mu_{\phi}(x),\,\operatorname{diag}\sigma_{\phi}^{2}(x)\big),$
@@
$KL\!\left(\mathcal{N}(\mu,\operatorname{diag}\sigma^{2})\,\middle\|\,\mathcal{N}(0,\sigma_{0}^{2}I)\right)=\tfrac{1}{2}\sum_{j}KL_{j},$
where $KL_{j}=(\sigma_{j}^{2}+\mu_{j}^{2})/\sigma_{0}^{2}-1-\log(\sigma_{j}^{2}/\sigma_{0}^{2})$. We sum this over all layers and parameters (both $W_{\ell}$ and $b_{\ell}$).
@@ 4 Experiments
We evaluate FDN on (i) controlled 1D regression tasks where the ground-truth function is known and interpolation vs. extrapolation is precisely defined in input space, and (ii) standard UCI-style regression benchmarks with feature-based ID/OOD splits. In both settings we report metrics separately on the interpolation (ID) region, the extrapolation (OOD) region, and the aggregated test set, and we summarize distribution shift by deltas $\Delta(\cdot)=\mathbb{E}_{\mathrm{OOD}}[\cdot]-\mathbb{E}_{\mathrm{ID}}[\cdot]$.
@@ 4.3 Complexity, capacity, and fairness
Link-budget. We consider networks with a single hidden layer and fix the parameter budget to $P\approx 1000$ (±5%) for all models, counting all trainable parameters, including any Hypernetwork components; counts appear in Table 4 (Appendix F). To equalize the update budget, ensembles with $M$ members use epoch-split training (epochs divided by $M$). For non-ensemble networks we use one Monte Carlo draw per update ($K=1$), keeping per-step cost comparable and the total number of parameter updates matched across models.
@@ 4.4 Metrics and calibration diagnostics
To avoid cherry-picking, all qualitative plots are based on a representative seed selected by a fixed aggregation rule. For each dataset and configuration, we run all models over multiple random seeds (20 for toy tasks, 100 for Airfoil, and 3 for the remaining UCI datasets) and compute per-seed summary metrics (e.g., ID/OOD MSE, variance, $\Delta\mathrm{CRPS}$, AURC). For IC-FDNet (or LP-FDNet), we take the coordinate-wise median of these metrics and select the seed whose metric vector is closest (Euclidean distance) to this median. All qualitative plots use this representative seed under the same rule across datasets, ensuring the shown behavior is typical rather than hand-picked. For Airfoil, Figure 5 (Appendix F) confirms that the selected seed lies close to the overall across-seed trend.
@@ 4.5 Results
Real regression benchmarks. On the real regression tasks (Airfoil, CCPP, Energy), the same patterns largely persist (Figures 3 and 4; Tables 810). FDN achieves reasonable in-distribution MSE and typically exhibits large positive $\Delta\mathrm{Var}$, indicating that uncertainty widens under feature-based shift. This comes with somewhat larger $\Delta\mathrm{MSE}$ and $\Delta\mathrm{CRPS}$ than the sharpest baselines on Airfoil, but more moderate values on CCPP and Energy, so calibration degrades gradually rather than catastrophically. On Airfoil, for example, FDN’s ID scatter lies close to the ideal MSE==Var line and spreads out smoothly OOD, with strong Spearman correlation; on CCPP and Energy the predictive variances remain shift-aware and provide useful selective-risk behavior, even when the absolute scale is not always better than the strongest baselines. Taken together, the toy and UCI results suggest that input-conditioned weight stochasticity is a viable and modular route to OOD-aware regression: FDN behaves like a drop-in “uncertainty layer” that can be tuned via $\beta$ and hypernetwork capacity to trade off in-distribution sharpness against OOD conservatism.
@@ 5 Limitations
FDN’s input-conditioned weight stochasticity, like other stochastic layer designs, can overfit spurious cues if $\beta$ is too small or the prior is too loose, making careful KL scheduling and prior choice important. LP-FDN samples weights layer-by-layer, introducing additional latency relative to a deterministic forward pass; test-time sampling also incurs a compute–latency trade-off with the number of Monte Carlo samples $K$, though we keep $K$ small in our experiments. Our study focuses on low-dimensional regression with tabular datasets and modest network backbones; extending FDN to high-dimensional inputs (e.g., images), deeper architectures, or structured outputs will likely require additional engineering, such as low-rank or adapter-style hypernetworks and tighter capacity control. On the calibration side, FDN can under-scale variance in highly oscillatory OOD regimes (e.g., the sine task), and the current design does not explicitly enforce frequency-aware or spectral robustness.